OpenStudy (anonymous):

Find P(A or B), if events A and B are overlapping. P(A) =9/18 P(B) =5/18 P(A and B) = 3/18

OpenStudy (amistre64):

p(a or b) = p(a) + p(b) - p(a and b)

OpenStudy (amistre64):

9+5 = 14 - 3 = 11 11/18

OpenStudy (amistre64):

anytime ther is a value for p(a and b) it means they overlap and has to be subtracted out for the double count

OpenStudy (anonymous):

Thank You Again!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (amistre64):

yeps :)